If 4 factories in a group of 20 factories in a community are violating environmental regulations and 6 are randomly selected for inspection, what is the probability that none of the violators will be selected?
I did: 6/20= 0.3% no violators will be selected
Is this correct?
Do I leave it as a percentage?

Answer by stanbon(57282)

If 4 factories in a group of 20 factories in a community are violating environmental regulations and 6 are randomly selected for inspection, what is the probability that none of the violators will be selected?
-----------
16 of the 20 factories are not violators.
P(not violator) = 16/20 = 4/5 = 0.8
P(violator) = 0.2
-----------------------
Binomial Problem with n= 6 and p = 0.8
P(x=6 non-violators) = 6C6(0.8)^6(0.2)^0 = 0.8^6 = 0.2621
=================
Cheers,
Stan H.